# Md Kowsher School of Engineering and Science

### v1

@@ Abstract
In Chaos, a minor divergence between two initial conditions exhibits exponential amplification over time, leading to far-away outcomes, known as the butterfly effect. Thus, the distant future is full of uncertainty and hard to forecast. We introduce Group Reservoir Transformer to predict long-term events more accurately and robustly by overcoming two challenges in Chaos: (1) the extensive historical sequences and (2) the sensitivity to initial conditions. A reservoir is attached to a Transformer to efficiently handle arbitrarily long historical lengths, with an extension of a group of reservoirs to reduce the uncertainty due to the initialization variations. Our architecture consistently outperforms state-of-the-art DNN models in multivariate time series, including NLinear, Pyformer, Informer, Autoformer, and the baseline Transformer, with an error reduction of up to -89.43% in various fields such as ETTh, ETTm, and air quality, demonstrating that an ensemble of butterfly learning, the prediction can be improved to a more adequate and certain one, despite of the traveling time to the unknown future.
@@ 1 Introduction
In the realm of addressing time-series tasks, prior research has explored a multitude of techniques, with a notable emphasis on the utilization of deep neural network models, exemplified by Transformer (Zeng et al., 2022). Nevertheless, an intrinsic constraint associated with the Transformer architecture, originally proposed by (Vaswani et al., 2017), manifests in its quadratic time and memory complexity concerning input length. This inherent limitation imposes a significant hindrance when endeavoring to perform long-term forecasting tasks. Several notable efforts, including UnlimiFormer (Bertsch et al., 2023), Efficient Transformer (Tay et al., 2022), and Reformer (Kitaev et al., 2020), have been made to extend the input length of Transformers. These endeavors primarily hinge on modifications to the attention model itself, often founded on certain assumptions. However, the outcome of these approaches is the extension of the input length to a fixed value, thereby yielding limited adaptability for learning from and predicting arbitrary long sequences.
We directly tackle the challenges associated with stable and accurate long-term chaotic prediction, aiming to push the boundaries of existing technology in time series forecasting. Our approach represents an innovative solution tailored to the complex task of extremely long-term chaotic prediction. We integrate reservoir computing into the Transformer framework, offering a novel approach that can effectively model sequences of inputs with arbitrary lengths. Reservoir computing, known for its simplicity and computational efficiency, excels in processing temporal data within the context of chaotic time series prediction (Bollt, 2021). Within our architecture, the reservoir is instrumental in capturing the intricate dependencies across all temporal events without window-sized context restriction; for instance, it discerns the sequence of events based on all historical data. Subsequently, the Transformer digests short-term event chronicles, melding its insights with those of the reservoir to produce the final forecast. This paradigm departures from conventional heuristic assumptions, ushering in a structured method that propels transformative changes in chaotic forecasting domain.
Our framework grapples with two technical obstacles related to Lyapunov time and the butterfly effect. Firstly, the confined input length of the Transformer considerably constrains the reservoir’s dimensions and potential. This limitation arises since the number of reservoir nodes scales quadratically with the reservoir output size, which mirrors the input size of DNN models, including the Transformer. To efficiently manage longer sequences, our system employs a nonlinear readout rooted in single-head attention mechanisms, superseding the traditional linear reservoir readout. Given its heightened expressiveness, this nonlinear readout essentially performs a dimensionality reduction on reservoir outputs, thus providing the Transformer with more meaningful and potent feature inputs.
@@
Empirical results demonstrate that the nonlinear readout, when sourced from ensemble reservoirs, drastically refines the Transformer’s prowess in time-series predictions, registering an error reduction of up to 89.43%. Through experiments, we gauge the chaotic nature inherent in our time-series datasets. Furthermore, our methodology surpasses leading-edge DNNs, aptly handling very long input lengths, signaling a transformative phase for the Transformer model.
Our contributions are mainly four folds: a) We integrate reservoir computing into the system, enabling the Transformer to tackle long-term predictions effectively. b) We refine traditional reservoir computing techniques by substituting the linear readout with a nonlinear counterpart, facilitating streamlined dimensionality reduction and adept feature assimilation. c) To address the nuances of reservoir initialization sensitivity, we implement ensemble reservoirs. d) Our approach demonstrated superior performance in forecasting chaotic phenomena, accurately predicting Lyapunov time steps across multiple horizons. Furthermore, it achieved state-of-the-art results in forecasting multivariate time series data.
@@ 2 Background
Consider a historical time series data containing $\vec{u}(t)$, a vector of features with $c$ components, and $\vec{y}(t)$ is corresponding labels for each time stamp $t\in T$, where $T$ is the total time stamps in a finite history. Now we aim to forecast $\vec{y}(t+1)$ the class or the value of a future time step given $t$ sequential multivariate history $\vec{u}(1:t+1)$ and $\vec{y}(1:t)$.
Time-Series Forecasting often only handles fixed input length with an assumption of looking back $s$ window size and considering $(t-s)$ history to predict at $t$ instead of using the complete history. Then the task is to predict $\vec{y}(t+1)$ given $\vec{u}(t-s:t+1),\vec{y}(t-s:t)$ with the conditional distribution:
$Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t);\phi)$ (1)
Here $\phi$ indicates learnable parameters shared by all-time steps $T$ to predict the conditional probability. The reason to introduce $s$ is that learning this conditional probability typically depends on $s$. For example, Transformer takes $\mathcal{O}(s^{2}\times c$). The computational cost is quadratic to the $s$.
We introduce deep reservoir computing for time series in chaotic prediction to avoid this costly operation. The advantage of reservoir computing is to preserve all the history for prediction. Therefore, it becomes possible to estimate conditional distribution with any length input:
$Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t);\phi)$ (2)
If the reservoir has $m$ size of the output vector, the time complexity of the proposed work will be $\mathcal{O}((k+m)^{2}\times c)$. Here, the parameter $k$ represents the size of the small look-back window compared to $s$.
@@ 2.1 Deep Reservoir Computing
Our basic reservoir architecture follows the deep reservoir computing (DRC) (Gallicchio et al., 2017) with the Echo State Network (ESN) model and a trained linear readout. In the basic Leaky Integrator ESN (LI-ESN) model Gallicchio et al. (2017), the state is updated according to the following state transition function:
$\vec{x}(t)=(1-a)\vec{x}(t-1)+\alpha\tanh(\vec{W}_{in}\vec{u}(t)+\bm{\theta}+\vec{\hat{W}}\vec{x}(t-1)),$ (3)
where $\vec{x}(t)\in R^{n}$ denotes the reservoir state at time $t$, $n$ represents the dimensionality of the reservoir state vector. $\vec{W}_{in}\in R^{c\times n}$ is the input-to-reservoir weight matrix, $\bm{\theta}\in R^{n}$ is the bias-to-reservoir weight vector, $\vec{\hat{W}}\in R^{n\times n}$ is the recurrent reservoir weight matrix, $\tanh(.)$ is the element-wise applied hyperbolic tangent activation function, and $\alpha\in[0,1]$ is the leaky parameter, i.e., the decay rate of the nodes. The reservoir parameters are initialized according to the constraints specified by the Echo State Property (ESP) and then are left untrained Gallicchio & Micheli (2017). Accordingly, the weight values in $\vec{W}_{in}$ and in $\bm{\theta}$ are chosen from a uniform distribution over $[-{scale}_{in},{scale}_{in}]$, where ${scale}_{in}$ represents an input-scaling parameter. Matrix $\vec{\hat{W}}$ contains randomly selected values from a uniform distribution. It is then adjusted so that the spectral radius of matrix $\vec{\tilde{W}}=(1-\alpha)I+\alpha\vec{\hat{W}}$, denoted as $\rho$, remains below 1 ($\rho<1$). This process helps ensure stable dynamics in the reservoir system. The reservoir weight $\vec{\hat{W}}$ and input weight $\vec{W_{in}}$ are fixed and random, where each weight is drawn according to Gaussian distribution with parameterized variances.
@@ 2.2 Linear Readout
$y(t)=\vec{W}_{out}{\vec{x}(t)}^{T}+\theta_{out},$ (4)
where $y(t)\in R^{m}$ is the linear readout with $m$ dimensionality. $\vec{W}_{out}\in R^{n\times m}$ represents the reservoir-to-readout weight matrix, connecting the reservoir units to the units in the readout. Since only $\vec{W}_{out}$ are trained, the optimization problem boils down to linear regression. Training is typically not a limiting factor in DRC, in sharp contrast with other neural network architectures. The expressiveness and power of reservoir computing rather lie in the high-dimensional non-linear dynamics of the reservoir.
The reservoir is effective for managing lengthy inputs. When we combine it with a Transformer model, the time complexity becomes $\mathcal{O}((k+m)^{2}\times c)$. For the Transformer, if we include more time steps in the input, the training time increases quadratically with the input length. On the other hand, the reservoir treats each time step individually, keeping the entire history in its learning process without making the input longer. This makes the reservoir a better choice for handling long sequences of input data.
2.3 Baseline Transformer
In Transformer, the architecture unravels the temporal dependencies into individual time steps (Wu et al., 2020). At each time of training, we optimize $\phi$ parameters by a Transformer $\mathcal{M}$ to learn the distribution $Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t))$ composed of two inputs: the current time stamp $\vec{u}(t+1)$ and that of the previous $s$ time steps $\vec{u}(t-s:t),\vec{y}(t-s:t)$:
$\vec{g}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)(\vec{u}(t-s:t),\vec{y}(t-s:t))$ (5)
Here $\kappa$ is a learning parameter adjusting the weights of the current input and previous inputs. Thus $\vec{g}(t+1)$ is the input to Transformer, $\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{g}(t+1))$. Nonetheless, employing $\vec{g}(t+1)$ as the input for the Transformer poses two significant challenges. Firstly, the Transformer’s input length is constrained by a factorization approach involving a retrospective window of size $s$, thereby increasing computational complexity. Secondly, the Transformer imposes restrictions on input length, preventing it from exceeding the value of $s$ due to quadratic time complexity considerations. Consequently, the incorporation of extensive historical data becomes impractical, limiting its potential impact on the learning and decision-making processes.
3 Method
3.1 Readout with Self-Attention
One downside of the reservoir is that if the reservoir output size $m$ is large, the Transformer training is slow due to its quadratic complexity. Since a linear readout of a reservoir needs a much larger size to have the same expressive power as the nonlinear readout, we introduce nonlinear readout layers, which work quite well in reducing DRC output dimension and improving prediction performance. Replacing linear readout with nonlinear readout leads to faster convergence and reduces the possibility of getting stuck in a local optimum (Triefenbach & Martens, 2011).
To model nonlinear readout, we use attention mechanisms (Vaswani et al., 2017) to capture input feature importance, alleviating the problem of vanishing gradient of long-distance dependency. There have been various implementations of attention mechanisms. We implement our attention model as in (Zhao, 2022). Attention mechanisms process sequential data that considers the context for each timestamp. $W$ and $b$ are this representation’s weight and bias, and the hyperbolic tangent function $\tanh(.)$ is a nonlinear activation.
$h_{i,j}=\tanh(\vec{x}(t)_{i}^{T}\vec{W_{t}}+\vec{x}(t)_{j}^{T}\vec{W_{x}}+b_{i})$ (6)
$e_{i,j}=\sigma(\vec{W_{\alpha}}h_{i,j}+b_{\alpha})$ (7)
$\gamma_{i,j}=\frac{exp(e_{i,j})}{\sum_{i=1}^{J}(exp(e_{i,j}))}$ (8)
$r(t)=\left\{\sum_{j}\gamma_{i,j}\vec{x}(t)_{j}\right\}_{i=1}^{c}$ (9)
Here, $r(t)$ is a non-linear readout using the attention mechanism and the subscript $i$ and $j$ are the $i$-th and $j$-th components of $\vec{x}(t)$. The attention weight $\gamma$ is a softmax of $e$.
@@ 3.2 Group Reservoir
In chaos theory, the butterfly effect is the sensitive dependence on initial conditions in which a small change in one state of a deterministic nonlinear system can result in large differences in a later state. To enhance reservoir performance, we integrate multiple reservoirs’ nonlinear readout layers, ensuring their independence. We consider $Q$ reservoirs with distinct decay rates ($\alpha$ and $\rho$) initialized randomly Gallicchio et al. (2017). Utilizing Equations 3 and 9, if $r^{l}(t)$ represents the nonlinear readout from the attention mechanism of the $l$-th reservoir within $Q$, we combine various leaky parameters ($\alpha$) and spectral radius ($\rho$) across the reservoirs. This results $\vec{o}(t)$ in a grouped ESN representation in element-wise $\oplus$ addition to all reservoirs’ attention outputs:
$\vec{o}(t)=r^{1}(t)\oplus r^{2}(t)\oplus...\oplus r^{Q}(t)$ (10)
3.3 Reservoir Transformer (RT)

### v2

# Changes by Butterflies: Farsighted Forecasting with Group Reservoir Transformer

@@ Abstract
In Chaos, a minor divergence between two initial conditions exhibits exponential amplification over time, leading to far-away outcomes, known as the butterfly effect. Thus, the distant future is full of uncertainty and hard to forecast. We introduce Group Reservoir Transformer to predict long-term events more accurately and robustly by overcoming two challenges in Chaos: (1) the extensive historical sequences and (2) the sensitivity to initial conditions. A reservoir is attached to a Transformer to efficiently handle arbitrarily long historical lengths, with an extension of a group of reservoirs to reduce the sensitivity to the initialization variations. Our architecture consistently outperforms state-of-the-art models in multivariate time series, including TimeLLM, GPT2TS, PatchTST, DLinear, TimeNet, and the baseline Transformer, with an error reduction of up to -59% in various fields such as ETTh, ETTm, and air quality, demonstrating that an ensemble of butterfly learning can improve the adequacy and certainty of event prediction, despite of the traveling time to the unknown future.
@@ 1 Introduction
Prior research has explored a multitude of techniques in chaotic event prediction, with a notable emphasis on the utilization of deep neural network models, exemplified by Transformer (Zeng et al., 2022). Nevertheless, the Transformer architecture’s quadratic time and memory complexity of the input length significantly limits the performance of long-term forecasting tasks (Vaswani et al., 2017). Several notable efforts, including but not limited to UnlimiFormer (Bertsch et al., 2023), Efficient Transformer (Tay et al., 2022), and Reformer (Kitaev et al., 2020), extend the input length of Transformers but primarily hinge on modifications to the attention model itself, often founded on certain assumptions to extend the input length to a fixed value without considering the temporal dependency of the very far away events, thereby yielding limited adaptability for learning from and predicting arbitrary long sequences.
Stream reservoir learning: We integrate reservoir computing, as an autoregressive model, into the Transformer framework that effectively models sequences of inputs with arbitrary lengths. Reservoir computing, known for its simplicity and computational efficiency, excels in processing temporal data within the context of chaotic time series prediction (Bollt, 2021). Within our architecture, the reservoir captures dependencies across all temporal events by discerning the sequence of events based on all historical data without window-sized context restriction. The reservoir learning processes the input in a stream fashion and requires no training by freezing reservoir weights. Its linear time complexity and the constant space complexity to the input length make handling the arbitrary context length possible. Subsequently, the Transformer digests short-term events in quadratic time, melding its insights with long-term events modeled by the reservoir to produce a more accurate final forecast. Our framework grapples with two technical obstacles related to Lyapunov time and the butterfly effect using the following two techniques: nonlinear readout and group reservoir.
Non-linearity: The traditional linear readout is expensive since the number of reservoir nodes scales quadratically with the reservoir output size, which mirrors the input size of the Transformer. Our system employs a nonlinear readout rooted in single-head attention mechanisms to manage longer sequences efficiently, superseding the traditional linear reservoir readout. Given its heightened expressiveness, this nonlinear readout reduces reservoir outputs’ dimensionality, thus providing the Transformer with more meaningful and potent feature inputs and dimension reduction.
@@
Empirical results demonstrate that the nonlinear readout and group reservoir drastically improve the Transformer and its variants in time-series predictions with an error reduction of up to 59.70%. Our methodology surpasses Transformer and other leading-edge deep neural networks (DNNs). It aptly handles very long input lengths, signaling a notable improvement for the Transformer model.
Our contributions are mainly fourfold: (1) We integrate reservoir computing into the system, enabling the Transformer to tackle long-term predictions effectively. (2) We refine traditional reservoir computing techniques by improving the linear readout with a nonlinear counterpart, facilitating streamlined dimensionality reduction while keep the autoregressive learning on temporal patterns. (3) We implement the group reservoir to address the nuances of reservoir initialization sensitivity. (4) Our approach demonstrated superior performance in forecasting chaotic phenomena in long-term horizons. Furthermore, it achieved state-of-the-art results in forecasting multivariate time series data.
@@ 2 Background
Preliminary: We consider a time series dataset denoted by $\mathbf{u}_{1:T}$, where each element $\mathbf{u}_{t}\in\mathbb{R}^{N_{u}}$ represents a $N_{u}$-dimensional vector at time $t\in\{1,\cdots,T\}$. Given a rolling forecasting scenario with a fixed window size $w$, our objective is to predict $\tau$ future time steps, $\mathbf{u}_{t+1:t+\tau}$, based on the input $\mathbf{u}_{t-w:t}$. We propose the following model:
$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{u}_{t-w:t},\mathbf{\Phi})$ (1)
Here, $\mathbf{\Phi}$ represents learnable parameters, and $\mathbb{M}(\cdot)$ is a Transformer time-series forecasting model.
3 Method
Transformer has a quadratic time complexity in terms of the history length and prediction future time steps (Jin et al., 2024; Nie et al., 2022; Beltagy et al., 2020). To reduce the computational complexity of window sizes, we introduce reservoir computing $\mathcal{R}$ to model the entire history of the time series efficiently:
$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{u}_{t-k:t}\uplus\mathcal{R}(\mathbf{u}_{1:T}),\mathbf{\Phi})$ (2)
[…]
3.1 Cross-attention for Embedding
We first embed the time series data for the neural networks. Since the embedding window has a fixed size, time series sequential input usually are chunked into segments, and the learning happens segment by segment. To make each segment comparable to the others, the full time series input are often normalized. Here we use the reversible instance normalization technique for normalization (Kim et al., 2021; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021), where the full sequence $\mathbf{u}_{1:T}$ is normalized to have zero mean and unit variance. We denote each token embedding as $\epsilon(\mathbf{u_{t}})\in\mathbb{R}^{d_{\epsilon}}$, where $\quad\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$, and the input $\mathbf{u}_{1:T}$ is embedded into ${\epsilon({\mathbf{u}}_{1}),\epsilon({\mathbf{u}}_{2}),\ldots,\epsilon({\mathbf{u}}_{T}})$. To address the dependency of each time step in one segment, we apply the cross-attention (Zhang and Yan, 2022; Rombach et al., 2022), where each vector $\epsilon(\mathbf{u}_{t-k:t})$ considers the local temporal relationship with its $i$-window-sized neighbors $\epsilon(\mathbf{u}_{t-i:t+i})$:
$\mathbf{h}_{t}=\text{softmax}\left(\frac{\mathbf{W}_{k}\cdot\epsilon(\mathbf{u}_{t-k:t})(\mathbf{W}_{q}\cdot\epsilon(\mathbf{u}_{t-i:t+i}))^{T}}{\sqrt{d_{\epsilon}}}\right)\mathbf{W}_{v}\cdot\epsilon(\mathbf{u}_{t-i:t+i}),$ (3)
where $\mathbf{h}_{t}\in\mathbb{R}^{k\times d_{\epsilon}}$ carries the local temporal relationship for time step $t$. Here $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ are trainable matrices to learn queries, keys, and values, respectively (Wei et al., 2020).
@@ 3.2 Deep Reservoir Computing
The Echo State Network (ESN) (Jaeger and Haas, 2004) implements discrete-time dynamical systems as one of the reservoir computing frameworks (Lukoševičius and Jaeger, 2009; Verstraeten et al., 2007). The recurrent reservoir layers are untrained and provide a suffix-based Markovian representation of the past input history (Tiňo et al., 2007), attaching a trained linear readout.
We use the Leaky Integrator ESN (LI-ESN) model (Jaeger et al., 2007), a variant of the basic ESN, in which leaky integrator reservoir units are adopted. In a LI-ESN with $N_{u}$ input units and $N_{r}$ reservoir units the state is updated according to the following state transition function:
$\mathbf{x}_{t}=(1-\alpha)\mathbf{x}_{t-1}+\alpha\tanh(\mathbf{W}_{in}\mathbf{h}_{t}+\boldsymbol{\theta}+\mathbf{W}\mathbf{x}_{t-1}),$ (4)
where $\mathbf{x}_{t}\in\mathbb{R}^{N_{r}}$ denote the input and the reservoir state at time $t$, respectively. $\mathbf{W}_{in}\in\mathbb{R}^{N_{r}\times d_{\epsilon}}$ is the input-to-reservoir weight matrix, $\boldsymbol{\theta}\in\mathbb{R}^{N_{r}}$ is the bias-to-reservoir weight vector. $\mathbf{W}\in\mathbb{R}^{N_{r}\times N_{r}}$ is the recurrent reservoir weight matrix, $\tanh$ is the element-wise hyperbolic tangent activation function, and $\alpha\in[0,1]$ is the leaky parameter. The reservoir parameters are initialized according to the constraints specified by the Echo State Property (ESP) (Jaeger and Haas, 2004; Tiňo et al., 2007) and then are left untrained, which makes its computing efficient. Accordingly, the weight values in $\mathbf{W}_{in}$ and in $\boldsymbol{\theta}$ are chosen from a uniform distribution over $[-\sigma_{in},\sigma_{in}]$, where $\sigma_{in}$ represents an input-scaling parameter. The values in matrix $\mathbf{W}$ are randomly selected from a uniform distribution and then re-scaled such that the spectral radius (i.e. the largest eigenvalue in absolute value) of matrix $\mathbf{W}=(1-\alpha)I+\alpha\mathbf{W}$, denoted by $\rho$, is smaller than 1, i.e. $\rho<1$ (Jaeger and Haas, 2004). Note that the leaky parameter $\alpha$ and the spectral radius of the recurrent weight matrix $\rho$ represent two relevant reservoir computing hyper-parameters. $\alpha$ is set to a fixed value of $0.7$ for simplicity, and $\rho$ is uniform randomly sampled between [00,$1$].
@@ 3.3 Linear Readout
$\mathbf{y}_{t}=\mathbf{W}_{out}{\mathbf{x}_{t}}+\boldsymbol{\theta}_{out},$ (5)
where $\mathbf{y}_{t}\in\mathbb{R}^{m}$ is the linear readout with $m$ dimensionality. $\mathbf{W}_{out}\in\mathbb{R}^{m\times N_{r}}$ represents the reservoir-to-readout weight matrix, connecting the reservoir units to the units in the readout. $\boldsymbol{\theta}_{out}\in\mathbb{R}^{m}$ is the bias-to-readout weight vector. Training the readout involves solving a linear regression problem, typically approached by using Moore-Penrose pseudo-inversion, with further details in  Lukoševičius and Jaeger (2009); Jaeger and Haas (2004). Since only $\mathbf{W}_{out}$ are trained, the optimization problem boils down to linear regression.
3.4 Nonlinear Readout
The output of the reservoir is the input of the Transformer. One downside of the reservoir is that if the reservoir output size is large, then the Transformer training will be slow due to its quadratic complexity to its input size. Since a linear readout of a reservoir needs a much larger size to have the same expressive power as the nonlinear readout, we introduce nonlinear readout layers, which work quite well in reducing reservoir output dimension and improving prediction performance hereafter. Therefore, for a faster convergence and avoiding a local optimum (Triefenbach and Martens, 2011), we append a nonlinear readout to the linear readout using the following non-linear activation function:
$\mathbf{\Lambda}_{t}^{l}=\sigma(\mathbf{y}_{t}^{l}),$ (6)
where $\sigma(\mathbf{y}_{t}^{l})$ is the $l$-th reservoir output, and we use $L$ many reserviors in the following section.
@@ 3.5 Group Reservoir
In Chaos theory, the butterfly effect refers to the sensitive dependence on initial conditions, where a small change in one state of a deterministic nonlinear system can result in large differences in a later state. As ensemble reduces the variants of the prediction, we integrate multiple reservoirs’ nonlinear readout layers for the stability of our prediction. We consider $L$ reservoirs, each with distinct decay rates ($\alpha$ and $\rho$), initialized randomly (Gallicchio et al., 2017). This results in a grouped Echo State Network (ESN) representation $\mathbf{o}_{t}$, achieved by element-wise $\oplus$ addition of all reservoirs’ non-linear activation outputs:
$\mathbf{o}_{t}=\mathbf{\Lambda}_{t}^{1}\oplus\mathbf{\Lambda}_{t}^{2}\oplus\ldots\oplus\mathbf{\Lambda}_{t}^{L}$ (7)
3.6 Group Reservoir Readout with Self-Attention
In order to capture the input feature importance, alleviating the problem of vanishing gradient of long-distance dependency, we have used a self-attention mechanism from the aggregate output of group reservoir in Equation 7 by following Shaw et al. (2018) :
$e_{ij}=\frac{(\mathbf{o}_{t}^{i}W^{Q})(\mathbf{o}_{t}^{j}W^{K})^{T}}{\sqrt{d_{\epsilon}}}$ (8)
$\alpha_{ij}=\frac{\exp{e_{ij}}}{\sum_{k=1}^{n}\exp{e_{ik}}}$ (9)
$\mathbf{z_{t}}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{o}_{t}^{j}W^{V})$ (10)
3.7 Group Reservoir Transformer (RT)